Consider the following spot interest rates for maturities of one, two, three, and four years. r1=2.93% r2=3.71% r3=4.27% r4=3.74% What is the three year forward rate one year from now (in percent)? Use the exact formula. Answer to two decimals, carry intermediate calcs. to four decimals.
Answer #1

3 year forward rate one year from now = (1+r4)4/(1+r1) -1 = (1+3.74%)4/(1+2.93%) -1 = 12.52%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well